Find the sum of n term of A.P. whose rth term is 2r+3

Grade:11

2 Answers

Meet
6 years ago
The general term is 2r+3 so the sum is 2[n(n+1)/2] + 3n and it`s value is equal to n^2+4n..............
Aryan Vartak
6 years ago
r th term is 2r + 3So 2r + 3 = a + (r - 1)d 2r + 3 = a + dr - d(dr = 2r)(d = 2) 2r + 3 = a + 2r - 2 3 = a - 2(a = 5)Sum of n terms in AP is (n/2)[2a+(n-1)d]Substituting the values of `a` and `d`We get:(n/2)[(2*5)+(n-1)*2]All the 2`s in the equation cancel out, leaving:n(5+n-1) = n(4+n) = n^2 + 4n